OpenStudy (anonymous):

please help. if A={5 2} and B={0 -3}, what is 2 +3? -1 3 -2 1 A B

OpenStudy (amistre64):

multiply as required, and add like 'parts'

OpenStudy (amistre64):

2( 5,2) + 3( 0,-3) = 2(-1,3) + 3(-2, 1) =

OpenStudy (anonymous):

(10,-5) and (-6, 3) right?

OpenStudy (amistre64):

first one is good, try the other one (-6,3) doesnt work right

OpenStudy (anonymous):

(-8,9)????

OpenStudy (amistre64):

better yes :)

OpenStudy (anonymous):

okay what now?

OpenStudy (amistre64):

now input that as the answer of course ....

OpenStudy (amistre64):

2A + 3B = 2( 5,2) + 3( 0,-3) = (10 , -6) 2(-1,3) + 3(-2, 1) = (-8 , 9)

OpenStudy (amistre64):

10 , -5 , accursed typos

OpenStudy (anonymous):

the choices are a. [25 -5] b. [10 -5] c. [5 -1] d. [15 0] -15 20 -8 9 -3 4 -7 11

OpenStudy (amistre64):

so pick the one that matches .... wheres the disconnect here?

OpenStudy (anonymous):

oh.. b right?

OpenStudy (amistre64):

perfect :)

OpenStudy (anonymous):

ok thanks very much!

OpenStudy (amistre64):

good luck